# An extension of the Erdős-Turán additive base conjecture via generalized circles of partition

### 13 Feb 2024 v7

@@ 1. Introduction and background
It is worth observing that each axis is uniquely determined by points $[x]\in\mathcal{C}(n,\mathbb{M})$ and that each point of a CoP $\mathcal{C}(n,\mathbb{M})$ excluding an existing center has exactly one real axis partner. To see this, just observe that a degenerated axis is determined by the center of the CoP. And this is unique if it exists. Let $\mathbb{L}_{[x],[y]}$ be a real axis of the CoP $\mathcal{C}(n,\mathbb{M})$. Suppose as well that $\mathbb{L}_{[x],[z]}$ is also a real axis with $z\neq y$. It follows from the ensuing discussion that we must have $n=x+y=x+z$ and therefore $y=z$. This cannot be and the claim follows immediately.
@@ 2. Generalized circles of partition
Throughout this paper we will denote for simplicity the generalized circle of partition in simple wording as g-CoP. Also, it is worth pointing out that various basic features that holds for CoPs does hold for generalized CoPs, except for previously technical results that needs investigating. Next we introduce the notion of the concentration of g-CoPs.
@@ 2.1. Axial potential of generalized circles of partition
for any subset $\mathbb{A}\subset\mathbb{N}$. This does not hold if the underlying set $\mathbb{A}$ is highly ”irregular”. Indeed, under the requirement of the conjecture, that

### 08 Mar 2024 v8

@@ 1. Introduction and background
It is worth observing that each axis is uniquely determined by points $[x]\in\mathcal{C}(n,\mathbb{M})$ and that each point of a CoP $\mathcal{C}(n,\mathbb{M})$ excluding an existing center has exactly one real axis partner. To see this, just observe that a degenerated axis is determined by the center of the CoP and this is unique if it exists. Now, let $\mathbb{L}_{[x],[y]}$ be a real axis of the CoP $\mathcal{C}(n,\mathbb{M})$. Suppose as well that $\mathbb{L}_{[x],[z]}$ is also a real axis with $z\neq y$. It follows from the ensuing discussion that we must have $n=x+y=x+z$ and therefore $y=z$. This cannot be and the claim follows immediately.
@@ 2. Generalized circles of partition
Throughout this paper we will denote for simplicity the generalized circle of partition in simple wording as g-CoP. Also, it is worth pointing out that various basic features that holds for CoPs does hold for generalized CoPs, except for previously technical results that needs investigating. Next we introduce the notion of the axial potential of g-CoPs.
@@ 2.1. Axial potential of generalized circles of partition
Indeed there are many minor yet illuminating steps that have been omitted in this argument. Under the assumption of the conjecture, that
#$\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{A})\right\}>0$
for all sufficiently large values of $n$, then we can write
#$\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{A})\right\}=H(n)\lfloor\frac{|\mathbb{A}\cap\mathbb{N}_{n}|-1}{2}\rfloor\geq H(n)\lfloor\frac{n^{1-\epsilon}-1}{2}\rfloor$
for some $0<H(n)\leq 1$ for all sufficiently large $n$. Plugging this into the computation of the truncated $2^{nd}$ axial potential (for large $k$), we deduce
$\sum\limits_{n=3}^{k}\frac{\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{A},\mathbb{N})\right\}^{2}}{\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{N})\right\}^{2}}\geq\sum\limits_{n=3}^{k}\frac{\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{A})\right\}^{2}}{\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{N})\right\}^{2}}\geq\mathcal{P}(k)\sum\limits_{n=3}^{k}\frac{\lfloor\frac{n^{1-\epsilon}-1}{2}\rfloor^{2}}{\lfloor\frac{n-1}{2}\rfloor^{2}}$
for $\mathcal{P}(k)>0$. It follows that
$\sum\limits_{n=3}^{k}\frac{\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{A})\right\}^{2}}{\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(n,\mathbb{N})\right\}^{2}}\gg\mathcal{P}(k)\sum\limits_{n=3}^{k}\frac{1}{n^{2\epsilon}}.$
Clearly the right hand side, which is always a partial sum of positive terms for all large $k$, tends to infinity as $k\longrightarrow\infty$ for all $0<\epsilon\leq\frac{1}{2}$. If we assume on the contrary that $\lim\limits_{k\longrightarrow\infty}\mathcal{P}(k)\sum\limits_{n=3}^{k}\frac{1}{n^{2\epsilon}}<\infty$ for $0<\epsilon\leq\frac{1}{2}$ then
$0<\lim\limits_{k\longrightarrow\infty}\mathcal{P}(k)\sum\limits_{n=3}^{k}\frac{1}{n^{2\epsilon}}<\infty$
since the partial sums are the sum of positive terms for all sufficiently large $k$ under the assumption of the conjecture. It implies that
$\sum\limits_{n=3}^{\infty}\frac{1}{n^{2\epsilon}}<\infty$
for $0<\epsilon\leq\frac{1}{2}$ which is absurd.
@@
for any subset $\mathbb{A}\subset\mathbb{N}$. This does not hold if the underlying set $\mathbb{A}$ is highly ”irregular”. More precisely, if we take the set $\mathbb{A}$ to be the set of prime numbers $\mathbb{P}$ then we obtain
#$\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(30,\mathbb{P})\right\}=3$
since the weight set of the CoP is given by $||\mathcal{C}(30,\mathbb{P})||=\{7,11,13,17,19,23\}$ with the only axes $\mathbb{L}_{[7],[23]}$ and $\mathbb{L}_{[11],[19]}$ and $\mathbb{L}_{[13],[17]}$. However, we obtain
$\lfloor\frac{|\mathbb{P}\cap\mathbb{N}_{30}|-1}{2}\rfloor=4$
and we see that
#$\#\left\{\mathbb{L}_{[x],[y]}~{}\hat{\in}~{}\mathcal{C}(30,\mathbb{P})\right\}<\lfloor\frac{|\mathbb{P}\cap\mathbb{N}_{30}|-1}{2}\rfloor.$
The main point is that the first supposed inequality does not always hold for certain subset of the integers like the prime numbers. Nonetheless, it holds for CoPs with natural number base set. Indeed, under the requirement of the conjecture, that